OpenStudy (anonymous):

What is the 24th term of the arithmetic sequence 10,16,22,...?Answers to choose from. a)146 b)147 c)148 d)149

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

Firstly, find the common difference here...

OpenStudy (anonymous):

its 10+(24-1)*6=10+138=148

OpenStudy (anonymous):

Common Difference = Second Term - First Term..

OpenStudy (anonymous):

Tank you @matricked goodnight

OpenStudy (anonymous):

And just follow @matricked who has given you the answer without any hesitation..

OpenStudy (anonymous):

Everyone on openstudy should learn : http://openstudy.com/code-of-conduct So, must read this... Good Night @ridgeboy36 ..